# Composite Flow Matching for Reinforcement Learning with Shifted-Dynamics Data

### v3

@@ Abstract
Incorporating pre-collected offline data can significantly improve the sample efficiency of reinforcement learning (RL), but this benefit is often challenged by discrepancies between the transition dynamics of the offline data and the online environment. Existing methods typically address this issue by penalizing or filtering out offline transitions in high dynamics-gap regions. However, their estimation of the dynamics gap often relies on KL divergence or mutual information, which can be ill-defined when the source and target dynamics have different support. To overcome these limitations, we propose CompFlow, a method grounded in the theoretical connection between flow matching and optimal transport. Specifically, we model the online dynamics as a conditional flow built upon the output distribution of the offline flow, rather than learning it directly from a Gaussian prior. This composite structure offers two key advantages: (1) improved generalization for learning online dynamics, and (2) a principled estimation of the dynamics gap via the Wasserstein distance between offline and online transitions. Leveraging our principled estimation of the dynamics gap, we further introduce an optimistic active data collection strategy that prioritizes exploration in regions of high dynamics gap, and theoretically prove that it reduces the performance disparity with the optimal policy. Empirically, CompFlow outperforms strong baselines across various RL benchmarks with shifted dynamics.
@@ 1 Introduction
Reinforcement Learning (RL) has demonstrated remarkable performance in complex sequential decision-making tasks such as playing Go and Atari games [schrittwieser2020mastering, silver2017mastering], supported by access to large amounts of online interactions with the environment. However, in many real-world domains such as robotics [kober2013reinforcement, tang2025deep], healthcare [yu2021reinforcement], and wildlife conservation [xu2021dual, xu2021robust, kong2025robust, kong2025generative], access to such interactions is often prohibitively expensive, unsafe, or infeasible. The limited availability of interactions presents a major challenge for learning effective and reliable policies. To address this challenge and improve sample efficiency during online training, a promising strategy is to incorporate a pre-collected offline dataset generated by a previous policy [song2022hybrid, nakamoto2023cal]. This approach enables the agent to learn from a broader set of experiences, which can help accelerate learning and improve performance [song2022hybrid].
A critical challenge in RL with offline data arises when the transition dynamics of the offline dataset differ from those of the online environment, where the agent is actively interacting and learning [qu2025hybrid]. This issue, commonly referred to as shifted dynamics, can introduce severe mismatches that bias policy updates, destabilize the learning process, and ultimately degrade performance. For instance,In robotics, differences in physical parameters such as friction can result in transition dynamics that diverge from the source domain. For example, in the conservation domain, historical data collected from one region may not generalize well to another, as differences in environmental conditions and ecological patterns across regions can lead to variations in poaching behavior [xu2021dual].
To address these challenges, existing methods either penalize the rewards or value estimates of offline transitions with high dynamics gap [niu2022trust, lyu2024cross], or filter out such transitions entirely [wen2024contrastive]. However, these approaches face key limitations. Most notably, the estimation of the dynamics gap typically relies on KL divergence or mutual information, both of which can be ill-defined when the offline and online transition dynamics have different supports [arjovsky2017wasserstein, poole2019variational].
@@ 2.1 Problem Definition
Lemma 2.2 (Return Bound between Two Environments [lyu2024cross]).
@@
This bound highlights two key sources of return gap between domains: (1) the mismatch between the learned policy and the behavior policy in the offline dataset, and (2) the shift in environment dynamics. The former can be mitigated through behavior cloning [lyu2024cross, xu2023cross], while the latter can be addressed by filtering out source transitions with large dynamics gaps [wen2024contrastive, lyu2025cross]. A central challenge lies in accurately estimating this gap. Existing methods often rely on KL divergence or mutual information [niu2022trust, lyu2024cross], which can be ill-defined when the two dynamics have different supports.
@@ 2.2 Flow Matching
In this paper, we will adopt flow matching [lipman2023flow, liu2022flow, albergo2023building] to model transition dynamics in reinforcement learning, owing to its ability to capture complex distributions. Flow Matching (FM) offers a simpler alternative to denoising diffusion models [ho2020denoising, song2020score], which are typically formulated using stochastic differential equations (SDEs). In contrast, FM is based on deterministic ordinary differential equations (ODEs), providing advantages such faster inference, and often improved sample quality.
@@
where $x_{t}=(1-t)x_{0}+tx_{1}$ denotes the linear interpolation between $x_{0}$ and $x_{1}$. Using linear interpolation paths encourages the learned flow to follow nearly straight-line trajectories, which reduces discretization error and improves the computational efficiency of ODE solvers during sampling [liu2022flow].
@@ 2.3 Optimal Transport Flow Matching and Wasserstein Distance
In the training objective of Eq. 1, when sample pairs $(x_{0},x_{1})$ are drawn from the optimal coupling $q^{\star}$, the flow model trained with the linear path matching loss learns a vector field that approximates the optimal transport plan. The transport cost of the learned flow is equal to the Wasserstein distance
$\mathbb{E}_{x_{0}\sim p_{0}}\left[\left\|\psi_{\theta}(x_{0},1)-x_{0}\right\|_{2}^{2}\right]=W^{2}_{2}(p_{0},p_{1}).$
For theoretical justification, see Theorem 4.2 in [pooladian2023multisample].
In practice, the optimal coupling $\pi^{\star}$ is approximated using mini-batches by solving a discrete OT problem between empirical samples from $p_{0}$ and $p_{1}$. The use of mini-batch OT has also been shown to implicitly regularize the transport plan [fatras2019learning, fatras2021minibatch], as the stochasticity from independently sampled batches induces behavior similar to entropic regularization [cuturi2013sinkhorn]. Further details of the OT-FM training procedure are provided in Appendix C.
@@ 3.1 Estimating Dynamics Gap via Composite Flow
To address this, we follow prior works [kerrigan2024dynamic, generale2024conditional] that handle samples with distinct conditioning variables. The key idea is to incorporate the conditioning variables into the cost function when solving for the optimal transport coupling. Specifically, we define the cost between two sample pairs $(s_{\rm off},a_{\rm off},s^{\prime}_{\rm off})$ and $(s_{\rm on},a_{\rm on},s^{\prime}_{\rm on})$ as:
$c\big((s_{\rm off},a_{\rm off},s^{\prime}_{\rm off},s_{\rm on},a_{\rm on},s^{\prime}_{\rm on}\big)=\|s^{\prime}_{\rm off}-s^{\prime}_{\rm on}\|_{2}^{2}+\eta\left(\|s_{\rm off}-s_{\rm on}\|_{2}^{2}+\|a_{\rm off}-a_{\rm on}\|_{2}^{2}\right),$
where $\eta>0$ is a weighting coefficient that controls the influence of the conditional terms.
Then the training objective of the online flow follows
$\mathcal{L}_{\rm on}(\phi)=\mathbb{E}_{((s_{\rm off},a_{\rm off},s^{\prime}_{\rm off}\doteq x_{1}),(s_{\rm on},a_{\rm on},s^{\prime}_{\rm on}\doteq x_{2}))\sim q^{*},t\sim\mathcal{U}[1,2]}\left[\left\|v_{\phi}\left(x_{t},t,s_{\rm on},a_{\rm on}\right)-(s^{\prime}_{\rm on}-s^{\prime}_{\rm off})\right\|^{2}\right],$
where $q^{*}$ is the optimal coupling of empirical distribution $(s_{\rm off},a_{\rm off},s^{\prime}_{\rm off})$ generated by learned source flow and empirical transition $(s_{\rm on},a_{\rm on},s^{\prime}_{\rm on})$ from target domain replay buffer. $x_{t}$ is the linear interpolation between $x_{1}$ and $x_{2}$.
The complete training algorithms of the offline flow and the online flow are in Appendix C.
Proposition 3.3 (Informal; shared-latent coupling is $W_{2}$-optimal).
Given $(s,a)\in\mathcal{S}\times\mathcal{A}$. As $\eta\to\infty$ and the training batch size $\to\infty$, the Wasserstein distance between the source and target transition distributions satisfies
$W_{2}^{2}(p_{\rm off}(\cdot|s,a),p_{\rm on}(\cdot|s,a))=\mathbb{E}_{x_{0}\sim\mathcal{N}(0,\mathbf{I})}\left[\left\|\psi^{\text{off}}_{\theta}(x_{0},1|s,a)-\psi^{\text{on}}_{\phi}(\psi^{\text{off}}_{\theta}(x_{0},1|s,a),2|s,a)\right\|_{2}^{2}\right].$
Monte Carlo Estimator. This result yields a practical Monte Carlo estimator of the dynamics gap $\Delta(s,a)$ using shared latent variables:
$\widehat{\Delta}(s,a)=\left(\frac{1}{M}\sum_{j=1}^{M}\left\|\psi^{\text{off}}_{\theta}(x_{0}^{(j)},1|s,a)-\psi^{\text{on}}_{\phi}(\psi^{\text{off}}_{\theta}(x_{0}^{(j)},1|s,a),2|s,a)\right\|^{2}_{2}\right)^{\frac{1}{2}},\quad x_{0}^{(j)}\sim\mathcal{N}(0,\mathbf{I}).$ (3)
@@ 3.2 Data Collection at High Dynamics Gap Region
where $\beta$ is a hyperparameter that trades off return and exploration of underexplored dynamics. To ensure sufficient coverage during training, we can further incorporate stochasticity by adding small perturbations to the selected actions [fujimoto2018addressing] or following a stochastic policy [haarnoja2018soft] consistent with widely used deep RL frameworks.
@@ 3.3 Practical Implementation
The policy is updated using a combination of policy improvement and behavior cloning, with the following objective:
$\mathcal{L}_{\pi}=\mathbb{E}_{s\sim D_{\rm off}\cup\mathcal{D}_{\rm on},\,a\sim\pi_{\varphi}(\cdot|s)}[Q_{\varsigma}(s,a)]-\omega\,\mathbb{E}_{(s,a)\sim\mathcal{D}_{\rm off},\,\tilde{a}\sim\pi_{\varphi}(\cdot|s)}[\|a-\tilde{a}\|^{2}],$ (8)
where $\omega$ is a hyperparameter that balances policy improvement with imitation of source actions. The behavior cloning term encourages sampled actions from the policy to match those in the offline dataset as suggested in Lemma 2.2.
The pseudocode of CompFlow, instantiated with Soft Actor-Critic (SAC) [haarnoja2018soft], is presented in Appendix D.
@@ 4 Related Work
Online RL with offline dataset. Online RL often requires extensive environment interactions [silver2016mastering, ye2020towards], which can be costly or impractical in real-world settings. To improve sample efficiency, Offline-to-Online RL leverages pre-collected offline data to bootstrap online learning [nair2020awac]. A typical two-phase approach trains an initial policy offline, then fine-tunes it online [nair2020awac, lee2022offline, nakamoto2023cal]. However, conservative strategies used to mitigate distributional shift, such as pessimistic value estimation [kumar2020conservative], can result in suboptimal initial policies and limit effective exploration [mark2022fine, nakamoto2023cal]. To resolve this tension, recent methods propose ensemble-based pessimism [lee2022offline], value calibration [nakamoto2023cal], optimistic action selection [lee2022offline], and policy expansion [zhang2023policy]. Others directly incorporate offline data into the replay buffer of off-policy algorithms [ball2023efficient], improving stability via ensemble distillation and layer normalization. However, these works assume identical dynamics between the source and target environments. In contrast, our work explicitly addresses the dynamics shift between the offline and online environments.
@@
RL with dynamics shift. Our work is related to cross-domain RL, where the source and target domains share the same observation and action spaces but differ in transition dynamics. Prior work has addressed such discrepancies via system identification [clavera2018learning, du2021auto, chebotar2019closing], domain randomization [slaoui2019robust, tobin2017domain, mehta2020active], imitation learning [kim2020domain, hejna2020hierarchically], and meta-RL [nagabandi2018learning, raileanu2020fast], often assuming shared environment distributions [song2020provably] or requiring expert demonstrations. More recent work has relaxed these assumptions. One line of research focuses on reward modification, which adjusts the reward function to penalize source transitions that are unlikely under the target dynamics [eysenbach2020off, liu2022dara], or down-weights value estimates in regions with high dynamics gap [niu2022trust]. Another line of work explores data filtering, which selects only source transitions with low estimated dynamics gap [xu2023cross], using metrics such as transition probability ratios [eysenbach2020off], mutual information [wen2024contrastive], value inconsistency [xu2023cross], or representation-based KL divergence [lyu2024cross]. However, these metrics can become unstable or ill-defined when the transition dynamics have different supports, and value-based methods are prone to instability caused by bootstrapping bias.
@@
In contrast, our approach leverages the theoretical connection between optimal transport and flow matching to estimate the dynamics gap in a principled manner. A closely related work by [lyu2025cross] also applies optimal transport to quantify the dynamics gap. However, their setting assumes both the source and target domains are offline, and their method estimates the gap based on the concatenated tuple $(s,a,s^{\prime})$ observed in offline data, rather than comparing conditional transition distributions. As a result, their metric does not accurately capture the gap in transition dynamics and cannot be used to compute exploration bonuses, which require gap estimation conditioned on a given state-action pair.
@@
RL with diffusion and flow models. Diffusion models [ho2020denoising, song2020score, kong2024diffusion] and flow matching [lipman2023flow, esser2024scaling] have emerged as powerful generative tools capable of modeling complex, high-dimensional distributions. Their application in RL is consequently expanding. Researchers have employed these generative models for various tasks, including planning and trajectory synthesis [janner2022planning], representing expressive multimodal policies [chi2023diffusionpolicy, ren2025diffusion], providing behavior regularization [chen2024score], or augmenting training datasets with synthesized experiences. While these works often focus on policy learning or modeling dynamics within a single environment, our approach targets the transfer learning setting. Specifically, we address scenarios characterized by a dynamics gap between the offline data and the online environment. We utilize Flow Matching, leveraging its connection to optimal transport, to estimate this gap.

### v4

@@ Abstract
Incorporating pre-collected offline data can substantially improve the sample efficiency of reinforcement learning (RL), but its benefits can break down when the transition dynamics in the offline dataset differ from those encountered online. Existing approaches typically mitigate this issue by penalizing or filtering offline transitions in regions with large dynamics gap. However, their dynamics-gap estimators often rely on KL divergence or mutual information, which can be ill-defined when offline and online dynamics have mismatched support. To address this challenge, we propose CompFlow, a principled framework built on the theoretical connection between flow matching and optimal transport. Specifically, we model the online dynamics as a conditional flow built upon the output distribution of a pretrained offline flow, rather than learning it directly from a Gaussian prior. This composite structure provides two advantages: (1) improved generalization when learning online dynamics under limited interaction data, and (2) a well-defined and stable estimate of the dynamics gap via the Wasserstein distance between offline and online transitions. Building on this dynamics-gap estimator, we further develop an optimistic active data collection strategy that prioritizes exploration in high-gap regions, and show theoretically that it reduces the performance gap to the optimal policy. Empirically, CompFlow consistently outperforms strong baselines across a range of RL benchmarks with shifted-dynamics data.
@@ 1 Introduction
Reinforcement Learning (RL) has demonstrated remarkable performance in complex sequential decision-making tasks such as playing Go and Atari games [53, 55], supported by access to large amounts of online interactions with the environment. However, in many real-world domains such as robotics [26, 60], healthcare [69], and wildlife conservation [66, 67, 30, 29], access to such interactions is often prohibitively expensive, unsafe, or infeasible. The limited availability of interactions presents a major challenge for learning effective and reliable policies. To address this challenge and improve sample efficiency during online training, a promising strategy is to incorporate a pre-collected offline dataset generated by a previous policy [59, 44]. This approach enables the agent to learn from a broader set of experiences, which can help accelerate learning and improve performance [59].
A critical challenge in online RL with offline data arises when the transition dynamics in the offline dataset differ from those in the online environment where the agent actively interacts and learns [49]. This issue, commonly referred to as shifted dynamics, can create severe distribution mismatch, bias policy updates, destabilize learning, and ultimately degrade performance. For instance, in robotics, the transition dynamics specify how the robot’s state (e.g., position and velocity) evolves after executing an action. During deployment, changes in physical parameters such as surface friction can alter this state evolution, causing the true next-state distribution to deviate from that implied by the historical data. Similarly, in conservation planning, the transition dynamics describe how the spatial risk of poaching evolves in response to patrol actions. Data collected in one region may not transfer to another because differences in terrain, accessibility, and human activity patterns lead to different state transitions and behavioral responses under the same patrol strategy [66].
To address these challenges, existing methods either penalize the rewards or value estimates of offline transitions with high dynamics gap [46, 37], or filter out such transitions entirely [64]. However, these approaches face key limitations. Most notably, the estimation of the dynamics gap typically relies on KL divergence or mutual information, both of which can be ill-defined when the offline and online transition dynamics have different supports [2, 48].
@@ 2.1 Problem Definition
Lemma 2.2 (Return Bound between Two Environments [37]).
@@
This bound highlights two key sources of return gap between domains: (1) the mismatch between the learned policy and the behavior policy in the offline dataset, and (2) the shift in environment dynamics. The former can be mitigated through behavior cloning [37, 65], while the latter can be addressed by filtering out source transitions with large dynamics gaps [64, 39]. A central challenge lies in accurately estimating this gap. Existing methods often rely on KL divergence or mutual information [46, 37], which can be ill-defined when the two dynamics have different supports.
@@ 2.2 Flow Matching
In this paper, we will adopt flow matching [34, 36, 1] to model the transition dynamics, owing to its ability to capture complex distributions. Flow Matching (FM) offers a simpler alternative to denoising diffusion models [21, 57], which are typically formulated using stochastic differential equations (SDEs). In contrast, FM is based on deterministic ordinary differential equations (ODEs), providing advantages such faster inference, and often improved sample quality.
@@
where $x_{t}=(1-t)x_{0}+tx_{1}$ denotes the linear interpolation between $x_{0}$ and $x_{1}$. Using linear interpolation paths encourages the learned flow to follow nearly straight-line trajectories, which reduces discretization error and improves the computational efficiency of ODE solvers during sampling [36].
@@ 2.3 Optimal Transport Flow Matching and Wasserstein Distance
In the training objective of Eq. 1, when sample pairs $(x_{0},x_{1})$ are drawn from the optimal coupling $q^{\star}$, the flow model trained with the linear path matching loss learns a vector field that approximates the optimal transport plan. The transport cost of the learned flow approximates the Wasserstein distance
$\mathbb{E}_{x_{0}\sim p_{0}}\left[\left\|\psi_{\theta}(x_{0},1)-x_{0}\right\|_{2}^{2}\right]\approx W^{2}_{2}(p_{0},p_{1}).$
For theoretical justification, see Theorem 4.2 in [47].
In practice, the optimal coupling $\pi^{\star}$ is approximated using mini-batches by solving a discrete OT problem between empirical samples from $p_{0}$ and $p_{1}$. The use of mini-batch OT has also been shown to implicitly regularize the transport plan [13, 14], as the stochasticity from independently sampled batches induces behavior similar to entropic regularization [9]. Further details of the OT-FM training procedure are provided in Appendix C.
@@ 3.1 Estimating Dynamics Gap via Composite Flow
To address this issue, we follow prior works [23, 18] and incorporate the conditioning variables directly into the transport cost. Specifically, consider pairs of transition samples $(s_{\rm off},a_{\rm off},s^{\prime}_{\rm off})$ and $(s_{\rm on},a_{\rm on},s^{\prime}_{\rm on})$. We define the transport cost as
$c\big((s_{\rm off},a_{\rm off},s^{\prime}_{\rm off}),(s_{\rm on},a_{\rm on},s^{\prime}_{\rm on})\big)=\|s^{\prime}_{\rm off}-s^{\prime}_{\rm on}\|_{2}^{2}+\eta\Big(\|s_{\rm off}-s_{\rm on}\|_{2}^{2}+\|a_{\rm off}-a_{\rm on}\|_{2}^{2}\Big),$
where $\eta>0$ controls the strength of alignment between conditioning variables.
Let $q^{*}$ denote the optimal coupling between the empirical distributions of $(s_{\rm off},a_{\rm off},s^{\prime}_{\rm off})$ and $(s_{\rm on},a_{\rm on},s^{\prime}_{\rm on})$ under the above cost. The online flow is trained using the objective
$\mathcal{L}_{\rm on}(\phi)=\mathbb{E}_{((s_{\rm off},a_{\rm off},s^{\prime}_{\rm off}\doteq x_{1}),(s_{\rm on},a_{\rm on},s^{\prime}_{\rm on}\doteq x_{2}))\sim q^{*},t\sim\mathcal{U}[1,2]}\left[\left\|v_{\phi}\left(x_{t},t,s_{\rm on},a_{\rm on}\right)-(s^{\prime}_{\rm on}-s^{\prime}_{\rm off})\right\|^{2}\right],$ (3)
where $x_{t}$ is the linear interpolation between $x_{1}$ and $x_{2}$.
Proposition 3.3 (Informal; shared-conditioning coupling is $W_{2}$-optimal).
Assume that, when training the online flow in Eq. (3), the marginals over the conditioning variables are matched, i.e., $p_{\rm off}(s,a)=p_{\rm on}(s,a)$. Then, for any fixed $(s,a)\in\mathcal{S}\times\mathcal{A}$, as $\eta\to\infty$ and the minibatch size used to compute the OT coupling tends to infinity, the expected squared displacement induced by the composite flow recovers the squared 2-Wasserstein distance between the offline and online transition distributions:
$\mathbb{E}_{x_{0}\sim\mathcal{N}(0,\mathbf{I})}\left[\left\|\psi^{\rm off}_{\theta}(x_{0},1|s,a)-\psi^{\rm on}_{\phi}\!\left(\psi^{\rm off}_{\theta}(x_{0},1|s,a),2|s,a\right)\right\|_{2}^{2}\right]\;\longrightarrow\;W_{2}^{2}\!\left(p_{\rm off}(\cdot|s,a),p_{\rm on}(\cdot|s,a)\right).$
Monte Carlo Estimator. This result yields a practical Monte Carlo estimator of the dynamics gap $\Delta(s,a)$:
$\widehat{\Delta}(s,a)=\left(\frac{1}{M}\sum_{j=1}^{M}\left\|\psi^{\rm off}_{\theta}(x_{0}^{(j)},1|s,a)-\psi^{\rm on}_{\phi}(\psi^{\rm off}_{\theta}(x_{0}^{(j)},1|s,a),2|s,a)\right\|_{2}^{2}\right)^{1/2},\quad x_{0}^{(j)}\sim\mathcal{N}(0,\mathbf{I}).$ (4)
Remark 3.4 (Enforcing the shared marginal condition).
The marginal condition $p_{\rm off}(s,a)=p_{\rm on}(s,a)$ can be enforced by construction. In practice, we construct the empirical distribution of offline triples $(s_{\rm off},a_{\rm off},s^{\prime}_{\rm off})$ by first sampling $(s_{\rm off},a_{\rm off})$ from the online dataset $\mathcal{D}_{\rm on}$, and then generating $s^{\prime}_{\rm off}\sim p_{\rm off}(\cdot|s_{\rm off},a_{\rm off})$ using the pretrained offline flow. Separately, we sample online transitions $(s_{\rm on},a_{\rm on},s^{\prime}_{\rm on})$ from $\mathcal{D}_{\rm on}$. As a result, the two empirical measures used in the OT coupling have matched $(s,a)$-marginals in expectation, while their conditional transitions differ. The complete training algorithms of the offline flow and the online flow are in Appendix C.
@@ 3.2 Data Collection at High Dynamics Gap Region
where $\beta$ is a hyperparameter that trades off return and exploration of underexplored dynamics. To ensure sufficient coverage during training, we can further incorporate stochasticity by adding small perturbations to the selected actions [17] or following a stochastic policy [19] consistent with widely used deep RL frameworks.
@@ 3.3 Practical Implementation
The policy is updated by combining policy improvement with a behavior cloning regularizer, using the objective
$\mathcal{L}_{\pi}=\mathbb{E}_{s\sim\mathcal{D}_{\rm off}\cup\mathcal{D}_{\rm on},\,a\sim\pi_{\varphi}(\cdot\mid s)}\big[Q_{\varsigma}(s,a)\big]\;-\;\omega\,\mathbb{E}_{(s,a)\sim\mathcal{D}_{\rm off},\,\tilde{a}\sim\pi_{\varphi}(\cdot\mid s)}\big[\|a-\tilde{a}\|_{2}^{2}\big],$ (9)
where $\omega>0$ controls the trade-off between maximizing the estimated $Q$-value and staying close to the offline behavior. The behavior cloning term encourages the policy’s sampled action $\tilde{a}$ to match the offline action $a$, as motivated by Lemma 2.2.
The pseudocode of CompFlow, instantiated with Soft Actor-Critic (SAC) [19], is presented in Appendix D.
@@ 4 Related Work
Online RL with offline dataset. Online RL often requires extensive environment interactions [54, 68], which can be costly or impractical in real-world settings. To improve sample efficiency, Offline-to-Online RL leverages pre-collected offline data to bootstrap online learning [43]. A typical two-phase approach trains an initial policy offline, then fine-tunes it online [43, 32, 44]. However, conservative strategies used to mitigate distributional shift, such as pessimistic value estimation [31], can result in suboptimal initial policies and limit effective exploration [40, 44]. To resolve this tension, recent methods propose ensemble-based pessimism [32], value calibration [44], optimistic action selection [32], and policy expansion [70]. Others directly incorporate offline data into the replay buffer of off-policy algorithms [3], improving stability via ensemble distillation and layer normalization. However, these approaches typically assume that the offline dataset is generated under the same transition dynamics as the online environment. In contrast, our work explicitly accounts for dynamics shift between the offline data and the online environment.
@@
RL with dynamics shift. Our work is related to cross-domain RL, where the source and target domains share the same observation and action spaces but differ in transition dynamics. Prior work has addressed such discrepancies via system identification [8, 10, 5], domain randomization [56, 61, 41], imitation learning [24, 20], and meta-RL [42, 50], often assuming shared environment distributions [58] or requiring expert demonstrations. More recent work has relaxed these assumptions. One line of research focuses on reward modification, which adjusts the reward function to penalize source transitions that are unlikely under the target dynamics [12, 35], or down-weights value estimates in regions with high dynamics gap [46]. Another line of work explores data filtering, which selects only source transitions with low estimated dynamics gap [65], using metrics such as transition probability ratios [12], mutual information [64], value inconsistency [65], or representation-based KL divergence [37]. However, these metrics can become unstable or ill-defined when the transition dynamics have different supports, and value-based methods are prone to instability caused by bootstrapping bias.
@@
In contrast, our approach leverages the theoretical connection between optimal transport and flow matching to estimate the dynamics gap in a principled manner. A closely related work by [39] also applies optimal transport to quantify the dynamics gap. However, their setting assumes both the source and target domains are offline, and their method estimates the gap based on the concatenated tuple $(s,a,s^{\prime})$ observed in offline data, rather than comparing conditional transition distributions. As a result, their metric does not accurately capture the gap in transition dynamics and cannot be used to compute exploration bonuses, which require gap estimation conditioned on a given state-action pair.
@@
RL with diffusion and flow models. Diffusion models [21, 57, 27] and flow matching [34, 11] have emerged as powerful generative tools capable of modeling complex, high-dimensional distributions. Their application in RL is consequently expanding. Researchers have employed these generative models for various tasks, including planning and trajectory synthesis [22], representing expressive multimodal policies [7, 51], providing behavior regularization [6], or augmenting training datasets with synthesized experiences. While these works often focus on policy learning or modeling dynamics within a single environment, our approach targets the transfer learning setting. Specifically, we address scenarios characterized by a dynamics gap between the offline data and the online environment. We utilize Flow Matching, leveraging its connection to optimal transport, to estimate this gap.